OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9

OpenStudy (texaschic101):

4x - 5y = 9 --- subtract 4x from both sides -5y = -4x + 9 --- divide both sides by -5 y = 4/5y - 9/5

OpenStudy (texaschic101):

oops..typo y = 4/5x - 9/5

OpenStudy (anonymous):

Solution Steps 1. Multiply Terms 4xy*–5=9 2. Multiply Terms 4–5xy=9 3. Divide Both Sides by Term x=94–5y Answer: 94–5y

OpenStudy (anonymous):

omg -.-

OpenStudy (anonymous):

give me amedle

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

i will u one too

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

@ugh.math Your answer is incorrect.

OpenStudy (anonymous):

how? :o

OpenStudy (anonymous):

Solving for y means you want the y to be on one side of the equal sign and the other stuff on the other side. You want the y alone. So you want y = something.